Grade 11Thermal Physics

A bulb of 3 litre capacity filled with air is heated from 27°C to t°C. The air this expelled measures 1.45L at 17°C. Considering the pressure to be constant 1 atm throughout and ignoring the expansion of bulb, calculate t.

Pranav mohril
9 Years agoGrade 11

1 Answer

musafir
8 Years ago
\frac{P_{300}V_{300}}{n_{300}300}=\frac{P_{t}V_{t}}{n_{t}t}

P300 = Pt = 1 atm
V300 = 3L
Vt = 3L+x (x is expanded volume)
n300=n
so:
 \frac{3}{300}=\frac{3+x}{t}
 
Rearrange them 
 
t=\frac{300\times (3+x)}{3}----(1)
Given:
 
\frac{V_{290}}{T_{290}}=\frac{V_{t}}{t}

\frac{1.45}{290}=\frac{x}{t}
x=\frac{1.45\times t}{290}------(2)
Put the value of x from equation (2) to equation (1) and solve for t
t=600K